the perimeter of rectangle 240 cm. its length is decreased by 10% & its breadth increased by 20 we get the same perimeter find the length & breadth of rectangle
Answer by ankor@dixie-net.com(22740)

the perimeter of rectangle 240 cm. its length is decreased by 10% & its breadth increased by 20 we get the same perimeter find the length & breadth of rectangle
:
I don't think that this will be possible if you increase the width by 20, If
you mean 20% it becomes a do-able problem.
:
Original rectangle
2L + 2W = 240
Simplify, divide by 2:
L + W = 120
L = (120-W)
:
New rectangle
(L-.1L) + (W+.2W) = 120
.9L + 1.2W = 120
.9L + 1.2W = 120
.9L + 1.2W = 120
:
substitute (120-W for L
.9(120-W) + 1.2W = 120
108 - .9W + 1.2W = 120
1.2W - .9W = 120 - 108
.3W = 12
W = 12/.3
W = 40
:
L = 120 - 40 = 80 cm
;
:
Check by finding new rectangle dimension:
.9 * 80 = 72; 1.2 * 40 = 48
2(72) + 2(48) = 240